A pipe contains water at 5.6 × 105 Pa above
atmospheric pressure.
If you patch a 4.66 mm diameter hole in the
pipe with a piece of bubble gum, how much
force must the gum be able to withstand?
Answer in units of N.
1
Expert's answer
2011-01-25T11:20:47-0500
The force would be
F = P/S = P/(πd2/4) = 4* 5.6 × 105 Pa / (π*4.662× 10-4 m2) = 3.28 × 108 N.